four times the measure of an angle is 7 more than three times the measure of its supplement. what is the measure of the angle?
Found 2 solutions by nerdybill, JBarnum:
Answer by nerdybill(7384)

four times the measure of an angle is 7 more than three times the measure of its supplement. what is the measure of the angle?
.
Two angles are supplement if the sum is 180 degrees.
.
Let x = the angle
then
180-x = supplement of angle
.
4x = 3(180-x)+7
4x = 540-3x+7
7x = 540+7
7x = 547
x = 78.14 degrees

Answer by JBarnum(2146)

SUPPLEMENT MEANS 2 ANGLES THAT ADD TO 180, Find x.
4x=7+3y
x+y=180
y==180-x
4x=7+3(180-x)
4x=7+540-3x
7x=547